OpenStudy (anonymous):

What are the approximate solutions of 2x^2 -x + 10 = 0 ?

OpenStudy (triciaal):

one approach what are the factors of 20 such that when you add them you get -1?

OpenStudy (triciaal):

so using -5 and +4 2x^2 -5x +4x + 10 =0 regroup and factor 2x(x +2) -5(x +2) =0 common factor (x + 2) (2x-5)(x+2)=0 2x-5 = 0; x + 2 = 0 solve for x

OpenStudy (anonymous):

(2x - 5) (x + 2) . When you factor that out you get 2x^2 - x -10. It should be a positive 10. I believe you have to use the quadratic formula.

OpenStudy (triciaal):

you are right sorry but the method for factoring is correct the other approach is to use the quadratic formula first identify a, b and c using x = -b +/-rt of (b^2-4*a*c) / 2a plug the values of a, b and c then solve a = 2, b = -1 and c = 10